OpenStudy (anonymous):

Write the point-slope form of the equation passing through (5, -1) with a slope of 6. A. y-1=6(x+5) B. y+5=6(x-1) C. y+1=6(x-5) D. y-5=6(x+1) For this question all you have to do is plug n chug right? :D My answer was A. :)

OpenStudy (anonymous):

plug n chug works, but not very well. you should form the equation yourself. point slow are basically y-Y=m(x-X) where lowercase letters are variables and capitals are numbers so plug all your info in m is 6, and (5,-1) is your (X,Y) so y-(-1)=6(x-5) or y+1=6(x-5)

OpenStudy (anonymous):

oh okie dokie thanks :D I started to think maybe if I plugged them in the open that is equal to each other would work :) I was wrong always am :P

OpenStudy (anonymous):

that SHOULD work. BUT very often MORE THAN ONE equation will work for the same coordinates. the tests will try to trick you!!!

OpenStudy (anonymous):

Those evil test -.-

OpenStudy (anonymous):

haha one of my algebra classes just recently took a practice test where they had to simplify square roots. i knew exactly who plugged and who worked by the answers because every single answer was correct but only ONE was the MOST simplified XDDD

OpenStudy (anonymous):

so most kids chose random answers (since all the answers worked) but only a couple kids chose the correct answer

OpenStudy (anonymous):

Haha they always do find ways to make things more complicated than it should be....hey, it's like a bad relationship xD

OpenStudy (anonymous):

Yea, it's easy to tell from the people who did the work and people who just choose answers to get finished with it..

OpenStudy (anonymous):

thing is, math is extremely easy. its just they teach you boring crap first, then you develop this hatred for it, and soon really easy stuff looks really hard. imagine if you learned your ABC's for the first 12 years of school!! youd never learn to write and you would HATE it! thats what their doing with math! you're learning your "ABCs" for the first 12 years! by the time you graduate high school you only need to know how to use triangles... really?? no wonder people hate it! even the best books are horrible when read too slowly!

OpenStudy (anonymous):

wow that is like a spot on explanation :O That was 100% true!!

OpenStudy (anonymous):

haha, im assuming your relatively young (dont tell me your age), in highschool, correct?

OpenStudy (anonymous):

yes and im already hating it :)

OpenStudy (anonymous):

that info is confidential O_O

OpenStudy (anonymous):

lol i asked because i want you to do me a favor. do the following: grow up to be some famous brilliant person in some cool field. make lots of money change out education system so that perhaps your kids or grandkids wont live through the hell that we have

OpenStudy (anonymous):

I'll keep that in mind :)

OpenStudy (anonymous):

I'm going to be a vet though xD so not sure how I can change our school systems through that career...

OpenStudy (anonymous):

youd be surprised. some of the most famous mathematicians in history werent mathematicians. a few were gorgeous models, body builders, actors, and even *shudder* lawyers (curse those liberal arts people!)

OpenStudy (anonymous):

lol I'll try my best :D